Vostron Status

Our engineers use this page to keep you up-to-date with the latest status of our network and systems.

When an alert is raised at our Network Operations Centre, our systems will automatically update the page.

Current status

All systems and services are online

Previous status updates

25th August 2017 at 15:49: Resolved: Connectivity
6th July 2017 at 12:22: Resolved: Hosted Telephony
As of 11:41 all Telephony services were resumed and customers should not experience any more issues.

The cause of this is still under investigation, and a full report will be sent out in due-course.

Going forward, we are closely monitoring our Telephony platform and have engineers on stand-by should any complications re-occur.

If any problems persist, please restart/reboot/power cycle your handset or contact support on 02380 111 222.

We sincerely apologise for any inconvenience this disruption may have caused, and assure you that we will be doing our upmost to ensure that such issues are not experienced in the future.

Vostron Support Team
3rd July 2017 at 08:58: Completed: Hosted Telephony
The scheduled upgrade has been completed successfully - All systems are now operating normally.

Please send any questions, comments, or issues to support@vostron.com.

We thank you for your patience.

Kind regards,
Vostron Support Team
22nd June 2017 at 14:00: Resolved: Connectivity – Southampton Science Park
22nd December 2016 at 16:42: Completed: Connectivity – Southampton Science Park
29th July 2016 at 16:55: Resolved: Hosted Telephony
The fault is now fully resolved.
26th July 2016 at 11:21: Completed: Connectivity – Southampton Science Park
This work completed successfully with approximately 6 minutes downtime to customers at most.
20th July 2016 at 14:09: Resolved: Connectivity – Broadband (ADSL/VDSL)
We believe this incident is now by and large resolved.
7th June 2016 at 21:32: Resolved: Connectivity – Southampton Science Park
The Science Park contractors have completed their maintenance.
7th June 2016 at 10:55: Resolved: Connectivity
This appears to have been an isolated incident related to one of our upstream carriers.
6th June 2016 at 10:52: Resolved: Connectivity – Leased Lines
We have received an RFO from Colt as follows:

Colt experienced a major network incident which occurred on the 01st June and following a brief period of stability re-occurred on the 2nd of June in the London region. The major network incident was due to excessive testing by a customer of their broadcast services. The testing caused a spike in traffic utilisation in the network and event flooding. The customer has agreed to cease testing. Colt has mitigated further instances via revoking access control to the testing portal and reviewing all planned deliveries of new service. Colt is reviewing an upgrade plan to current capacity.
27th May 2016 at 18:22: Resolved: Connectivity – Southampton Science Park
We are classing this incident as resolved as we have seen no further issues today.
26th May 2016 at 06:56: Completed: Connectivity – Southampton Science Park
Works have now completed.

We are still seeing a handful of customers as offline. An engineer will be with the affected customers shortly.
12th May 2016 at 18:38: Completed: Management tools
This work is now complete.
11th May 2016 at 12:20: Completed: Connectivity – Southampton Science Park
Work is now complete.
28th April 2016 at 12:16: Completed: Management tools
This work has been cancelled.
6th April 2016 at 18:52: Completed: Co-location – Docklands West
Works have been completed and with no impact to any customer services.
4th April 2016 at 12:38: Resolved: General
Power has been stable for 30 minutes now and SSE engineers have left site. Our engineering team are now working to get the office resources fully functional and a normal service will resume shortly.
14th September 2015 at 08:40: Completed: Management tools
The works have been completed successfully.
10th September 2015 at 17:47: Resolved: Connectivity – Southampton Science Park
Our engineers have left site and have confirmed that all affected services are now running normally.

Status breakdown